OpenStudy (boxman61):

Write the complex number in standard form a+bi -2+3√-8 need process more than answer please

Directrix (directrix):

-2+3√-8 = -2 + 3 [√ (-1) * √ 8] = -2 + 3 [ i * 2√2] = -2 + (6 √ 2) i

OpenStudy (boxman61):

is that \[-2+6\sqrt{2i}\] ?

Directrix (directrix):

No, the i for the imaginary unit is free-standing and is not under the radical.

Directrix (directrix):